A swimming pool 6 meters by 10 meters long is surrounded by a walk of uniform width. The area taken up by the walk happens to equal the area of the pool. What is the width of the walk?
Answer by TimothyLamb(4379)

---
pool dims:
x = length
y = breadth
---
outer walk dims:
u = length
v = breadth
--
walk's uniform width:
w = width
---
outer walk dims relative to pool dims:
u = (x + w) = length
v = (y + w) = breadth
---
area of pool
p = xy
---
area of walk:
w = p
w = uv - p
p = uv - p
xy = uv - xy
2xy = uv
2xy = (x + w)(y + w)
2xy = xy + wx + wy + ww
xy = wx + wy + ww
ww + (x + y)w - xy = 0
ww + (10 + 6)w - (10)(6) = 0
ww + 16w - 60 = 0
---
the above quadratic equation is in standard form, with a=1, b=16, and c=-60
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 16 -60
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
w = 3.13552873
w = -19.1355287
---
negative width doesn't fit the problem statement, so use the positive root:
---
answer:
width of the walk = 3.13552873 m
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php